The density of nitrogen at S.T.P is1.251kgm^-3. calculate the r.m.s. velocity of nitrogen molecules.

1
Expert's answer
2022-02-15T16:17:37-0500

Density

"\rho=1.251kg/m^3"

STP

Pressure

"P=1.013\times10^{5}pa"

We know that

"v_{RMS}=\sqrt\frac{3P}{\rho}"

Put value


"v_{RMS}=\sqrt\frac{3\times1.013\times10^5}{1.251}=4.92\times10^2m/sec"